OpenStudy (bananas1234):

is this an inverse variation? p(x) = x/11

OpenStudy (tkhunny):

Do x and p increase together or does one go in the opposite direction from the other?

OpenStudy (bananas1234):

opposite i think?

jimthompson5910 (jim_thompson5910):

Hint: make a table using x values that are multiples of 11 use x = 0 x = 11 x = 22 x = 33 and so on

OpenStudy (julyahx1):

do u have a question u need help with?

OpenStudy (tkhunny):

"I think"? Why not prove it? Please demonstrate, perhaps using Mr. Thompson's suggestion. "I think"? No.

OpenStudy (mathmale):

You may find this problem easier if you'd re-write your "p(x) = x/11" as p(x)=(1/11)x, which has the form y=kx. Please rethink this problem from this new perspective.